# Sequential $\Upsilon(nS)$ suppression in high-multiplicity pppp collisions: multi-differential constraints on hadronic mechanisms

### 19 Apr 2026 v1

@@ 1 Introduction
What makes this conclusion more than an isolated quarkonium argument is the context in which it sits. The onset of $\Upsilon$ suppression in $pp$ lines up, in the same multiplicity range, with the strangeness enhancement seen by ALICE [13] — the oldest and cleanest QGP signal known from heavy-ion physics — and with the long-range two-particle ridge that developed into a unique feature of high-multiplicity $pp$ data at the LHC [17]. At still higher multiplicities, ALICE has recently reported the baryon–meson $v_{2}$ grouping [15], the mass-ordering switch that is the fingerprint of partonic, rather than hadronic, collectivity. Independently, the equation-of-state analysis of Campanini & Ferri [18] places a soft-to-hard transition band precisely where all these phenomena switch on. A Bjorken-type estimate of the energy density at the top of the accessible multiplicity range lands comfortably above the lattice-QCD deconfinement crossover [20, 21], yet modestly enough that the absence of jet quenching in the same collisions is not a contradiction but a natural consequence of the path-length and sensitivity scales involved. Seen together, these are a convergent set of observations that single out a specific class of mechanisms and close off every other.
@@ 2 Notation and observables
The paper is organised as follows. Sec. 3 presents the multiplicity dependence of $R_{21}$ and $R_{31}$ inclusive and at high $p_{T}$. Sec. 4 uses the cone-isolation test to falsify local hadronic mechanisms. Sec. 5 uses the azimuthal-sector data to remove any residual near-side scenario. Sec. 6 uses the transverse sphericity to falsify every mechanism that is a function of $N_{\mathrm{track}}$ alone. Sec. 7 introduces the LHCb non-prompt $\psi(2\mathrm{S})/J/\psi$ temporal constraint. Sec. 8 gathers the verdicts on the hadronic and string-based frameworks. Sec. 9 connects the onset to strangeness enhancement, to the ridge, and to the ALICE partonic-flow observation. Sec. 10 estimates the Bjorken energy density and connects it to the equation-of-state analysis of Campanini & Ferri. Sec. 11 explains why the absence of jet quenching in $pp$ is compatible with a partonic droplet.
@@ 4 Cone isolation constrains local hadronic CIM
In the Armesto–Gavin–Ferreiro comover model with Cornell-potential radii $r_{1S}=0.226$ fm, $r_{2S}=0.486$ fm, $r_{3S}=0.690$ fm, the geometric dissociation cross sections scale as $\sigma_{2S}/\sigma_{1S}\simeq 4.6$, $\sigma_{3S}/\sigma_{1S}\simeq 9.3$, and the survival probability is $P_{\mathrm{surv}}(nS)=\exp[-\sigma_{nS}\,\rho_{\mathrm{com}}\,L]$. If $\rho_{\mathrm{com}}$ is identified with the local hadron density around the $\Upsilon$ — the physically honest reading of the word “comover” — the dense cone must produce substantially stronger suppression than the empty cone.
Keeping the three levels separate: data: the empty-cone and dense-cone classes are statistically indistinguishable within uncertainty. Experimental constraint: whatever mechanism generates the suppression does not scale with the local track density around the $\Upsilon$; a local-density proxy is not the right variable. Interpretation: in its local-density form the CIM is in direct tension with the cone data; the framework can of course be preserved by promoting $\rho_{\mathrm{com}}$ to a global quantity (Sec. 6 will show that this move is itself in tension with the sphericity-resolved data), but its most natural, physically transparent reading does not survive the cone test. This closes the first door.
@@ 5 Azimuthal independence removes the “near-side” escape
The backward result is particularly instructive: backward tracks are geometrically far from the $\Upsilon$, so they cannot represent a local scattering environment. The fact that they correlate with the suppression at the same rate as forward tracks implies that what drives the suppression is a global activity variable, of which all three sector multiplicities are proxies. Azimuthal independence is therefore a necessary but not sufficient element: it rules out proximity-based mechanisms, but does not by itself rule out a mechanism in which $N_{\mathrm{track}}$ itself is the relevant variable. The next step discriminates this remaining possibility.
@@ 6 Sphericity constrains every $N_{\text{track}}$-only mechanism
Globally this corresponds to the two significances that can be read directly off the extreme classes shown in Fig. 6: the $R_{32}$-versus-$N_{\mathrm{track}}$ slope is compatible with zero in jet-like events ($S_{T}<0.55$, $1.9\sigma$) and reaches $5.0\sigma$ in isotropic events ($S_{T}>0.85$[1]. The sphericity filter therefore operates on both $R_{21}$ and $R_{31}$ and acts coherently on the differential ratio $R_{32}$.
@@
On the differential ratio $R_{32}$ it is important to keep the three levels separated. Data: at $N_{\mathrm{track}}\simeq 74$ the high-sphericity class has $R_{32}\simeq 0.86\times R_{32}(\mathrm{jet})$, a $14\%$ gap between topologies at the same $N_{\mathrm{track}}$ [1]. Experimental constraint: the sequential hierarchy is driven by the event topology at fixed multiplicity rather than being inherited from production kinematics, because production kinematics are fixed when $N_{\mathrm{track}}$ is fixed. Interpretation, with caveats: the $R_{32}$ evidence is most natural in a core–corona picture where the isotropic core preferentially dissolves the more loosely bound state, but a mixing between core and corona contributions, or a non-trivial differential feed-down pattern, could in principle shift the quantitative picture. We do not rely on a specific CIM prediction for $R_{32}$ here: the observed topology-resolved $14\%$ gap is a direct experimental fact, and its interpretation in terms of any specific model requires inputs (Cornell radii, $\rho_{\mathrm{com}}(N_{\mathrm{track}},S_{T})$, feed-down fractions) that are not uniquely fixed.
@@ 7 Temporal constraint: LHCb non-prompt $\psi(2\mathrm{S})/J/\psi$
Figure 8: LHCb $pp$ data at $\sqrt{s}=13$ TeV: normalised $\psi(2\mathrm{S})/J/\psi$ cross-section ratio as a function of the non-dimensionalised multiplicity, in five exclusive $p_{T}^{\psi}$ slices ($0.3<p_{T}<2$, $2<p_{T}<4$, $4<p_{T}<6$, $6<p_{T}<8$, $8<p_{T}<20$ GeV/$c$), separately for prompt (left) and non-prompt (right) production. Own plot from the numerical Tables 4–8 of Appendix A of Ref. [2]; the self-normalisation is performed slice by slice to the average of the five multiplicity bins of the same slice. For prompt production the low-$p_{T}$ slices show the strongest multiplicity dependence, and the high-$p_{T}$ slice ($8<p_{T}<20$, yellow crosses) becomes compatible with flat within uncertainty. For non-prompt production no $p_{T}$ slice shows a significant multiplicity trend. This is the $\psi(2\mathrm{S})/J/\psi$ analogue of the CMS $\Upsilon$ evidence in Fig. 2.
@@ 8 Review of hadronic and string-based frameworks
Dissociation by nearby hadrons; cross sections scale as $(r_{nS}/r_{1S})^{2}$, so the model predicts more suppression in the dense-cone class than in the empty-cone class. The CMS data show the two classes to be indistinguishable within uncertainty. Verdict: in direct tension with the cone test; the model in its local-density reading does not reproduce the sign of the predicted effect.
@@
Overlapping strings from many MPI form ropes with enhanced effective tension $\kappa_{\mathrm{eff}}>\kappa_{0}$, preferentially suppressing loosely-bound states. Cone: OK (string overlaps set at production time, not driven by which cone the fragments end up in). Azimuth: OK (strings approximately uniform in $\phi$). Sphericity: the direction of the effect is correct (isotropic events have more MPI and hence larger rope tension), but no published PYTHIA+rope simulation reproduces, at fixed parameters, the full jet-like (compatible with zero, $1.9\sigma$) vs. isotropic ($5.0\sigma$) gap observed by CMS. $p_{T}$-escape: the rope tension does not depend on $p_{T}^{\Upsilon}$, so the model has no native escape mechanism; the LHCb low-$p_{T}$ enhancement of suppression is not naturally reproduced. Verdict: qualitatively consistent with cone and azimuth, quantitatively unvalidated on sphericity, and structurally missing a native $p_{T}$-escape; not ruled out as a partial contribution but not a complete explanation on present evidence.
@@ 9 Multi-observable context: strangeness, ridge, partonic flow
10 Bjorken energy density: consistency check
As a consistency check, not as a pillar of the argument, we estimate the Bjorken energy density [19] reached at the top of the multiplicity range and compare it to the lattice-QCD crossover band. We stress upfront that in $pp$ every input of this estimate is model dependent: the formation time $\tau_{0}$, the transverse area $A_{T}$, and the midrapidity $dE_{T}/dy$ are all inferred indirectly. The numbers below should be read as an order-of-magnitude benchmark, not as a measurement.
$\varepsilon_{\mathrm{BJ}}=\frac{1}{\tau_{0}A_{T}}\,\frac{dE_{T}}{dy}\Big|_{y=0},\qquad\tau_{0}\approx 1~\mathrm{fm}/c,~A_{T}=\pi R_{\mathrm{HBT}}^{2}\approx 10.2~\mathrm{fm}^{2},$ (1)
with $R_{\mathrm{HBT}}\approx 1.8$ fm measured in high-multiplicity $pp$ [1] and
$\frac{dE_{T}}{dy}\Big|_{y=0}\approx\langle m_{T}\rangle\times\tfrac{3}{2}\times dN_{\mathrm{ch}}/d\eta\approx 0.45\,\mathrm{GeV}\times 1.5\times 26\approx 17.6~\mathrm{GeV},$ (2)
where $\langle m_{T}\rangle\approx 0.45$ GeV is the transverse-mass average over the $\pi$, K, p spectra measured by ALICE at $\sqrt{s}=7$ TeV [16] and the factor $3/2$ accounts for neutral particles in the isospin-symmetric limit. Substituting:
$\varepsilon_{\mathrm{BJ}}\approx 1.7~\mathrm{GeV/fm}^{3},\qquad T\approx\left(\frac{30\,\varepsilon_{\mathrm{BJ}}}{37\pi^{2}}\right)^{1/4}\approx 180~\mathrm{MeV},$ (3)
using the massless-QGP Stefan–Boltzmann relation. The translation from $\varepsilon_{\mathrm{BJ}}$ to $T$ is itself crude because it assumes thermalisation and an ideal-gas equation of state, neither of which is established in $pp$.
Consistency with lattice QCD and with Campanini & Ferri.
This estimate exceeds the lattice-QCD pseudo-critical density $\varepsilon_{\mathrm{pc}}\approx 0.5$ GeV/fm3 [20, 21] by a factor $\sim 3$, with a temperature of the same order as $T_{\mathrm{pc}}\approx 156$ MeV. On independent grounds, the equation-of-state analysis of Campanini & Ferri [18], performed on $pp$ and $p\bar{p}$ multiplicity-dependent observables, identifies a soft-to-hard transition band $dN_{\mathrm{ch}}/d\eta\approx 6$2424. This band overlaps with the density window in which the signals discussed above become active. We read the convergence not as proof of deconfinement but as a consistency check: at the multiplicities where $\Upsilon(n\mathrm{S})$ sequential suppression, strangeness enhancement and (at higher $dN_{\mathrm{ch}}/d\eta$) the baryon–meson $v_{2}$ grouping all switch on, the naive hadronic description reaches its plausible limit of applicability [22]. The central claims of the paper rest on the multi-differential CMS and LHCb data analysed in Secs. 47; the present section is offered only as a back-of-the-envelope indication that those data sit in a density regime where partonic degrees of freedom are not phenomenologically unreasonable.
@@ 11 Why jet quenching is not visible — and why that is consistent
Radiative parton energy loss in a dense medium scales as
$\Delta E\propto\hat{q}\,L^{2},$ (4)
with $\hat{q}$ the transport coefficient and $L$ the path length. In central PbPb the medium extends over $L_{\mathrm{PbPb}}\sim 5$66 fm; in $pp$ the droplet extends over $L_{pp}\sim 1$22 fm. At comparable $\hat{q}$:
$\frac{\Delta E_{pp}}{\Delta E_{\mathrm{PbPb}}}\sim\left(\frac{L_{pp}}{L_{\mathrm{PbPb}}}\right)^{2}\sim\frac{1}{25}\text{--}\frac{1}{36}.$ (5)
An effect that is tens of GeV in PbPb becomes a few hundred MeV in $pp$. Current jet measurements in $pp$ are not sensitive to losses of that magnitude: the hadronic calorimetry threshold for identifying quenching is $\mathcal{O}(\text{GeV})$. The effect is therefore below the detection floor, not absent.
@@
is structural. Had jet quenching been measured in $pp$ at $\varepsilon_{\mathrm{BJ}}\approx 1.7$ GeV/fm3, that would be in tension with our estimate and with the present interpretation. The experimental facts — strong sequential $\Upsilon$ suppression, strangeness enhancement, ridge, partonic $v_{2}$, no jet quenching — are all simultaneously consistent with a density regime in which $\Upsilon(n\mathrm{S})$ states, and not high-$p_{T}$ jets, are the uniquely sensitive probes.
@@ 12 Interpretation and closure
    Equation-of-state consistency: the Bjorken-type estimate of the energy density (understood as a consistency check, not a measurement) sits in the soft-to-hard transition band identified by Campanini & Ferri [18], above the lattice-QCD crossover [20, 21].

### 23 Apr 2026 v2

# Sequential Y(nS) suppression in high-multiplicity pp collisions: the experimental case for an early, globally correlated medium

@@ 1 Introduction
What makes this conclusion more than an isolated quarkonium argument is the context in which it sits. The onset of $\Upsilon$ suppression in $pp$ lines up, in the same multiplicity range, with the strangeness enhancement seen by ALICE [13] — the oldest and cleanest QGP signal known from heavy-ion physics — and with the long-range two-particle ridge that developed into a unique feature of high-multiplicity $pp$ data at the LHC [17]. At still higher multiplicities, ALICE has recently reported the baryon–meson $v_{2}$ grouping [15], the mass-ordering switch that is the fingerprint of partonic, rather than hadronic, collectivity. Independently, the equation-of-state analysis of Campanini & Ferri [18] places a soft-to-hard transition band in $pp$ and $p\bar{p}$ data across ISR-to-LHC energies precisely where all these phenomena switch on. The absence of jet quenching in the same high-multiplicity $pp$ collisions is not a contradiction: path-length scaling alone suppresses any quenching signal below current detector sensitivity for a droplet of $\mathcal{O}(1)$ fm. Seen together, these are a convergent set of experimental observations that single out a specific class of mechanisms.
@@ 2 Notation and observables
The paper is organised as follows. Sec. 3 presents the multiplicity dependence of $R_{21}$ and $R_{31}$ inclusive and at high $p_{T}$. Sec. 4 uses the cone-isolation test to falsify local hadronic mechanisms. Sec. 5 uses the azimuthal-sector data to remove any residual near-side scenario. Sec. 6 uses the transverse sphericity to falsify every mechanism that is a function of $N_{\mathrm{track}}$ alone. Sec. 7 introduces the LHCb non-prompt $\psi(2\mathrm{S})/J/\psi$ temporal constraint. Sec. 8 gathers the verdicts on the hadronic and string-based frameworks. Sec. 9 connects the onset to strangeness enhancement, to the ridge, and to the ALICE partonic-flow observation. Sec. 10 notes the coincidence of the onset multiplicity window with the transition band of Campanini & Ferri. Sec. 11 explains why the absence of jet quenching in $pp$ is compatible with a partonic droplet.
@@ 4 Cone isolation constrains local hadronic CIM
Any mechanism in which the quarkonium is dissociated by inelastic scattering with nearby produced hadrons requires the dissociation probability to grow monotonically with the local hadron density around the $\Upsilon$ [6, 7, 8]. This expectation is model-independent: the word “comover” itself implies that proximity matters, and more comovers in the cone must produce more suppression, irrespective of the precise form of the dissociation cross section or the binding potential. The sign of the effect is unambiguous.
Keeping the three levels separate: Data: the empty-cone and dense-cone classes are statistically indistinguishable within uncertainty across the full $N_{\mathrm{track}}$ range accessible to both. Experimental constraint: the suppression does not scale with the local track density around the $\Upsilon$; whatever variable drives the suppression, it is not the number of hadrons collinear with the quarkonium. Interpretation: in its local-density form, any comover-like model is in direct tension with the cone data. The framework can of course be preserved by promoting the comover density to a global quantity; Sec. 6 shows that this move is itself in tension with the sphericity-resolved data. This closes the first door.
@@ 5 Azimuthal independence removes the “near-side” escape
The backward result is particularly instructive: backward tracks are geometrically far from the $\Upsilon$, so they cannot represent a local scattering environment. Yet the most compact and complete statement is provided by the transverse sector ($\pi/3<|\Delta\phi|<2\pi/3$): tracks at 90° to the $\Upsilon$ flight direction are neither co-moving nor recoiling against it, and have no preferred geometric or kinematic relationship to the $\Upsilon$ trajectory. Their multiplicity nonetheless correlates with the suppression at the same rate as forward tracks. The three-fold azimuthal symmetry — forward, transverse, and backward sectors all yielding statistically identical suppression slopes — is therefore a stronger statement than a forward–backward comparison alone: it closes not only the near-side escape but also any model based on recoil, deflection, or directional energy flow. What drives the suppression is a global activity variable, of which all three sector multiplicities are equivalent proxies. Azimuthal independence is therefore a necessary but not sufficient element: it rules out proximity-based and direction-based mechanisms, but does not by itself rule out a mechanism in which $N_{\mathrm{track}}$ itself is the relevant variable. The next step discriminates this remaining possibility.
@@ 6 Sphericity constrains every $N_{\text{track}}$-only mechanism
The two extreme classes show a striking contrast. In jet-like events ($S_{T}<0.55$) the slope of $R_{21}$ versus $N_{\mathrm{track}}$ is compatible with zero ($0.2\sigma$) and that of $R_{31}$ is marginal ($2.4\sigma$): the suppression is essentially absent. In isotropic events ($S_{T}>0.85$) the slope of $R_{21}$ differs from zero by $6.0\sigma$ and that of $R_{31}$ by $10.0\sigma$. The larger significance for $R_{31}$ reflects both its steeper slope (the more loosely bound state is more strongly suppressed) and the smaller relative uncertainties on $R_{31}$ in the high-$N_{\mathrm{track}}$ isotropic bins. Significance values are computed here from weighted linear fits to the seven $N_{\mathrm{track}}$ bins of the public HEPData tables of Ref. [1]; the null hypothesis is a slope of zero. The two populations have the same $N_{\mathrm{track}}$ but different topology; the suppression therefore depends on topology, not on multiplicity alone.
@@
Keeping the three levels separate: Data: at the highest accessible $N_{\mathrm{track}}$ the isotropic class shows $R_{21}$ suppressed by ${\sim}15\%$ and $R_{31}$ by ${\sim}25\%$ relative to the jet-like class at the same $N_{\mathrm{track}}$ [1]; the effect is stronger for the more loosely bound state. Experimental constraint: the suppression depends on the topology of the event at fixed multiplicity, not on multiplicity alone; and the sequential hierarchy — $R_{31}$ more suppressed than $R_{21}$ — is amplified in the isotropic class relative to the jet-like one. Interpretation, with caveats: the topology dependence is most naturally read as evidence that the effective medium density, rather than the track count, is what drives the dissociation; but a non-trivial differential feed-down pattern between the two event classes could in principle shift the quantitative picture.
@@ 7 Temporal constraint: LHCb non-prompt $\psi(2\mathrm{S})/J/\psi$
Figure 8: LHCb pppp data at $\sqrt{s}=13$ TeV: normalised $\psi(2\mathrm{S})/J/\psi$ cross-section ratio as a function of the non-dimensionalised multiplicity, in five exclusive $p_{T}^{\psi}$ slices ($0.3<p_{T}<2$, $2<p_{T}<4$, $4<p_{T}<6$, $6<p_{T}<8$, $8<p_{T}<20$ GeV/$c$), separately for prompt (left) and non-prompt (right) production. Own plot from the numerical Tables 4–8 of Appendix A of Ref. [2]; the self-normalisation is performed slice by slice to the average of the five multiplicity bins of the same slice. For prompt production the low-$p_{T}$ slices show the strongest multiplicity dependence, and the high-$p_{T}$ slice ($8<p_{T}<20$, yellow crosses) becomes compatible with flat within uncertainty. For non-prompt production no $p_{T}$ slice shows a significant multiplicity trend. This is the $\psi(2\mathrm{S})/J/\psi$ analogue of the CMS $\Upsilon$ evidence in Fig. 2.
@@ 8 Review of hadronic and string-based frameworks
Dissociation by nearby hadrons; the model predicts more suppression in the dense-cone class than in the empty-cone class. The CMS data show the two classes to be indistinguishable within uncertainty. Verdict: in direct tension with the cone test; the model in its local-density reading does not reproduce the sign of the predicted effect.
@@
Overlapping strings from many MPI form ropes with enhanced effective tension $\kappa_{\mathrm{eff}}>\kappa_{0}$, preferentially suppressing loosely-bound states. Cone: OK (string overlaps set at production time, not driven by which cone the fragments end up in). Azimuth: OK (strings approximately uniform in $\phi$). Sphericity: the direction of the effect is correct (isotropic events have more MPI and hence larger rope tension), but no published PYTHIA+rope simulation reproduces, at fixed parameters, the observed gap between the two extreme classes: $R_{21}$ compatible with flat ($0.2\sigma$) and $R_{31}$ marginal ($2.4\sigma$) in jet-like events, against $6.0\sigma$ and $10.0\sigma$ respectively in isotropic events at the same $N_{\mathrm{track}}$ [1]. $p_{T}$-escape: the rope tension does not depend on $p_{T}^{\Upsilon}$, so the model has no native escape mechanism; the LHCb low-$p_{T}$ enhancement of suppression is not naturally reproduced. Verdict: qualitatively consistent with cone and azimuth, quantitatively unvalidated on sphericity, and structurally missing a native $p_{T}$-escape; not ruled out as a partial contribution but not a complete explanation on present evidence.
@@ 9 Multi-observable context: strangeness, ridge, partonic flow
10 Onset multiplicity window: consistency with the Campanini & Ferri EOS
The four differential constraints discussed in Secs. 47 all become active in the same multiplicity window: $dN_{\mathrm{ch}}/d\eta\approx 4$66 at midrapidity. This is the onset of the suppression signal, of the strangeness enhancement (Sec. 9), and of the ridge; the baryon–meson $v_{2}$ grouping switches on at $dN_{\mathrm{ch}}/d\eta\gtrsim 17$.
On independent grounds, the equation-of-state analysis of Campanini & Ferri [18], performed on inclusive $pp$ and $p\bar{p}$ multiplicity distributions across ISR-to-LHC energies, identifies a soft-to-hard transition band at $dN_{\mathrm{ch}}/d\eta\approx 6$2424. This band is determined from the EOS proxy $\sigma_{s}/\langle p_{T}\rangle^{3}$ without reference to any quarkonium observable. The coincidence of the two onset windows is noted here as a consistency observation. It suggests that the density regime in which the multi-differential CMS and LHCb constraints point toward an early, globally correlated medium is also the regime in which the simplest hadronic description reaches the limit of its applicability. No energy-density estimate is required to reach this conclusion: the common variable throughout is $dN_{\mathrm{ch}}/d\eta$, a directly measured, model-independent quantity.
@@ 11 Why jet quenching is not visible — and why that is consistent
The path-length dependence of parton energy loss depends on the dominant mechanism. For radiative (gluon-radiation) loss in the LPM regime,
$\Delta E_{\mathrm{rad}}\propto\hat{q}\,L^{2},$ (1)
with $\hat{q}$ the transport coefficient and $L$ the path length. For purely collisional (elastic) loss the scaling is linear, $\Delta E_{\mathrm{coll}}\propto\hat{q}^{\prime}\,L$. In central PbPb the medium extends over $L_{\mathrm{PbPb}}\sim 5$66 fm; in $pp$ the droplet extends over $L_{pp}\sim 1$22 fm. The two parametrisations bracket the uncertainty:
$\frac{\Delta E_{pp}}{\Delta E_{\mathrm{PbPb}}}\sim\begin{cases}\left(\dfrac{L_{pp}}{L_{\mathrm{PbPb}}}\right)^{2}\sim\tfrac{1}{25}\text{--}\tfrac{1}{36}&\text{(radiative)}\\[8.0pt] \dfrac{L_{pp}}{L_{\mathrm{PbPb}}}\sim\tfrac{1}{4}\text{--}\tfrac{1}{5}&\text{(collisional)}\end{cases}$ (2)
An effect that is tens of GeV in central PbPb becomes at most a few GeV even in the most favourable (linear) scaling, and more likely a few hundred MeV in the radiative regime. In both cases it falls below the current sensitivity of jet measurements in $pp$, which require $\mathcal{O}(\text{GeV})$-scale modifications to be resolved above the hadronic calorimetry threshold. The effect is therefore below the detection floor, not absent, under any plausible scaling assumption.
@@
is structural. The experimental facts — strong sequential $\Upsilon$ suppression, strangeness enhancement, ridge, partonic $v_{2}$, no jet quenching — are all simultaneously consistent with a density regime in which $\Upsilon(n\mathrm{S})$ states, and not high-$p_{T}$ jets, are the uniquely sensitive probes.
@@ 12 Interpretation and closure
    Onset window coincidence: the multiplicity range in which all four differential constraints become active ($dN_{\mathrm{ch}}/d\eta\approx 4$66) coincides with the soft-to-hard transition band identified by Campanini & Ferri [18] from inclusive $pp$ and $p\bar{p}$ data across ISR-to-LHC energies, without reference to any quarkonium observable.
@@
Use of AI assistance
The author used Claude (Anthropic, claude-sonnet-4-6) for assistance with LaTeX editing, numerical cross-checks, and formal writing tasks during the preparation of this manuscript. All scientific ideas, analyses, interpretations, and conclusions are the author’s own and have been critically reviewed by the author.